An artist wants her painting to be in the shape of a golden rectangle. If the length of the painting is 36 inches, then what should be the width?
Answer by stanbon(75887)

An artist wants her painting to be in the shape of a golden rectangle. If the length of the painting is 36 inches, then what should be the width?
The golden ratio is 1/1.618
--------
Let width be "x" inches.
x/36 = 1/1.168
x = 36/1.168
x = 30.8219 inches (this could be the width)
---------------
OR
x/36 = 1.168/1
x = 36*1.168
x = 42.048 inches (this could be the width)
====================
Cheers,
Stan H.